# Towards Interpretability Without Sacrifice: Faithful Dense Layer Decomposition with Mixture of Decoders

### v1

@@ A.1 Proof of rank equality
Equation 6 shows that for a fixed expert $n$, the $n^{\text{th}}$ row $\mathbf{c}_{n}\in\mathbb{R}^{O}$ essentially scales the columns of matrix $\mathbf{D}\in\mathbb{R}^{H\times O}$. This is equivalent to right-multiplying matrix $\mathbf{D}$ by a diagonal matrix formed from $\mathbf{c}_{n}\in\mathbb{R}^{O}$. Indeed, the $(h,o)$ entry of such matrix product is
@@
since all off-diagonal terms (i.e., $i\neq o$) in Equation 7 vanish and $\text{diag}(\mathbf{c}_{n})_{oo}=c_{no}$ by construction. Comparing Equation 6 and Equation 9 shows that, for every $h\in\{1,2,\ldots,H\}$ and $o\in\{1,2,\ldots,O\}$ we have
@@
Finally, a standard result in linear algebra [90] has that $\text{rank}(\mathbf{AB})=\text{rank}(\mathbf{A})$ for any $\mathbf{A}\in\mathbb{R}^{H\times O}$ and invertible matrix $\mathbf{B}\in\mathbb{R}^{O\times O}$. Since matrix $\mathrm{diag}\left({\mathbf{c}}_{n}\right)\in\mathbb{R}^{O\times O}$ is invertible by assumption in Lemma 1, setting $\mathbf{A}=\mathbf{D}$ and $\mathbf{B}=\text{diag}({\mathbf{c}}_{n})$ yields the rank equality. ∎
@@ A.2 Proof of MxD forward pass equivalence
Recall we have input vector $\mathbf{z}\in\mathbb{R}^{H}$, expert coefficients $\mathbf{a}\in\mathbb{R}^{N}$, and layer weights $\boldsymbol{\mathcal{W}}\in\mathbb{R}^{N\times H\times O}$. The weights are defined in Equation 4 element-wise through the Hadamard product * as
@@
for learnable parameters $\mathbf{C}\in\mathbb{R}^{N\times O},\,\mathbf{D}\in\mathbb{R}^{H\times O}$. Lemma 2 states that MxD’s forward pass can be equivalently expressed as
@@
Plugging in the definition of $\mathbf{w}_{nh:}\in\mathbb{R}^{O}$ from Equation 4 then yields
@@
which is exactly the RHS of Equation 5, showing the MxD forward pass is equivalent to the Hadamard product of $\mathbf{C}^{\top}\mathbf{a}$ and $\mathbf{D}^{\top}\mathbf{z}$. ∎
@@ A.3 Intuition for weight parameterization through the lens of tensor methods
A second complementary way of viewing the MxD layer’s parameterization (and its full-rank properties) is through the lens of tensor methods [34]. A tensor-based motivation for MxD’s weight tensor parameterization and forward pass is presented in Section A.3.1 and Section A.3.2, respectively.
@@ A.3.1 MxD weight tensors through the Khatri-Rao product
MxDs construct the collective weight tensor through the Khatri-Rao product $\odot$ [34] of the two factor matrices $\mathbf{C}\in\mathbb{R}^{N\times O},\,\mathbf{D}\in\mathbb{R}^{H\times O}$. Concretely, the mode-$3$ unfolding222which is simply a reshaping of a higher-order tensor into a matrix, arranging all $N$ expert matrices’ column vectors along the columns of a new matrix. of the third-order weight tensor $\boldsymbol{\mathcal{W}}\in\mathbb{R}^{N\times H\times O}$ in MxDs from Equation 4 is alternatively given by:
@@ A.3.2 Tensorized MxD forward pass
Furthermore, the layer’s forward pass can then be viewed as performing two tensor contractions between the third-order weight tensor $\boldsymbol{\mathcal{W}}\in\mathbb{R}^{N\times H\times O}$ (collecting all $N$ experts’ $H\times O$-dimensional matrices) and expert coefficients $\mathbf{a}\in\mathbb{R}^{N}$ and hidden activations $\mathbf{z}\in\mathbb{R}^{H}$. This can be expressed in terms of the so-called mode-$n$ product (denoted by ×$\times_{n}$) [34] as follows:
@@ A.4 GLU encoders are a mixture of rank-1 linear experts
for input vector $\mathbf{x}\in\mathbb{R}^{I}$, learnable weights $\mathbf{E}_{\text{GLU}},\,\mathbf{E}\in\mathbb{R}^{I\times H}$, and activation function $\psi(.)$. To transform Equation 17 into the same model form as MxDs, we first pre-multiply the LHS by the identity matrix to match the MxD model form of Equation 5, yielding:
@@
where $(\mathbb{I})_{n}\in\mathbb{R}^{H}$ is the $n^{\text{th}}$ row of the $H$-dimensional identity matrix (i.e. a one-hot vector with its only non-zero element at index $n$). We draw particular attention to how the $n^{\text{th}}$ expert’s matrix $\mathbf{W}_{n}=\mathbf{E}\,\mathrm{diag}\left((\mathbb{I})_{n}\right)\in\mathbb{R}^{I\times H}$ essentially picks out the $n^{\text{th}}$ column of $\mathbf{E}$, leaving all remaining $H-1$ columns as zero vectors. Therefore, GLU encoders compute a MoE with linear expert weights of (at most) rank 1. This relationship between GLUs and conditional computation is consistent with prior work interpreting individual GLU column vectors as experts [92]. Whilst GLUs’ encoders’ model form does not put any inherent restrictions on the total number of rank-1 terms that can contribute to the output, the sparsity necessary for specialization does.
@@
    Replace $\mathbb{I}$ in Equation 18 with learnable, non-diagonal weight matrices for full-rankness.
@@ A.5 Hadamard-factorized tensors generalize MoVs
Prior work [37] proposes to linearly combine $N$ many (IA)3 adapters [38] for parameter-efficient MoEs for instruction fine-tuning. The implementation results in a very similar functional form to the factorized forward-pass in MxDs. Interestingly, the Hadamard product parameterization of the third-order weight tensor in Equation 4 provides a more general framework through which one can also derive MoVs’ model form, shedding light on the relationship to the proposed MxDs and their benefits. Concretely, factorizing the weight tensor instead along the second mode as $\boldsymbol{\mathcal{W}}(n,:,o)=\mathbf{c}_{n}*\mathbf{d}_{o}\in\mathbb{R}^{H}$ in our framework immediately recovers MoV [37] as a special case. In particular, in contrast to the MxD in Section A.3 whose weight tensor can be parametrized equivalently through its mode-$3$ unfolding [34], MoV’s implicit weight tensor can be given in terms of its mode-$2$ unfolding in terms of a similar Khatri-Rao product of two factor matrices.
Instead, MoVs in analogy would yield expert weights by pre-multiplying $\mathbf{D}$ as: $\mathbf{W}_{n}=\mathrm{diag}(\mathbf{c}_{n})\,\mathbf{D}\in\mathbb{R}^{H\times O}$ for much larger $\mathbf{C}\in\mathbb{R}^{N\times H}$). Due to $H\gg O$, our proposed MxD formulation yields around 4×4$\times$4 × the number of specialized units as MoVs with the same parameter budget (yet MoVs’ experts are of no higher rank than MxDs’), making MxDs a much more suitable and efficient class of layer for our goal of scalable specialization. We therefore see that the proposed lens of tensor methods for unification provides valuable insights about how to design more interpretable layers with the minimum trade-off to capabilities.
@@ B.1 Faithfulness in output space
Our main experiments measure model faithfulness in latent space–how well the sparse layer variants reconstruct the intermediate MLPs’ mapping. Here, we provide additional experiments comparing the faithfulness of sparse layers as their computation propagates to the model output space. Concretely, we sample $32$ consecutive tokens with the base model and then measure how similar the same generations are when the target MLP layer is replaced with the sparse layers.
We sample $512$ text snippets from OpenWebText, and use the first 4 words of each as the initial prompts, generating $32$ future tokens after each prompt. We plot in Figures 6 and 7 the percentage of the samples’ continuations that are identical in the original LLM and hooked LLMs up to $n$ future tokens ahead. We note that this is a rather punishing task–any small deviations quickly compound as $n$ grows. Despite this, we see that the MxDs match the future token generations far better than the baselines, exhibiting more faithfulness in model output space (as well as in latent space).
We also show qualitative examples of the first $8$ prompts and the subsequent ‘diffs’ (using Python 3’s difflib) of the generated tokens in Figures 8 and 9, where MxDs’ superior preservation can be viewed qualitatively.
Figure 6: Pythia-410m: The percentage of $512$ generated samples that contain $n$ words identical to the original model’s output (when replacing the base LLM’s MLP layer with the sparse layers).
Figure 7: GPT2-124m: The percentage of $512$ generated samples that contain $n$ words identical to the original model’s output (when replacing the base LLM’s MLP layer with the sparse layers).
@@ B.2 Additional reconstruction metrics
To highlight the scale of difference in the reconstructions between MxDs and the baselines, we also plot in Figure 10 the normalized MSE at the end of training for all models and LLMs. At the smallest values of $K$ (which we care about most for interpretability), MxDs’ normalized MSE is up to an order of magnitude smaller than Transcoders’.
@@ B.3 Results on additional layers
We also fully train all models and baselines (with 4 different values of $K$) on different target layers for each model. The results are shown in Figure 11 for $48$ additional trained layers for the same setup in the original paper, using different colours to highlight that these are new results. As can be seen, the same trend holds: MxDs significantly outperform the baselines at small $K$ in all LLMs.
@@
B.4 Expert rank
This section concerns the matrix rank of the linear experts in parameter-efficient MoEs. We first compare to low-rank MoEs in Section B.4.1 to demonstrate the benefits of full-rankness, and then follow up in Section B.4.2 by confirming that the learned MxD expert ranks are close to maximum in the trained models.
B.4.1 Comparisons to low-rank MoEs
In this section, we study the impact of expert rank on the ability of efficient MoE layers to reconstruct pre-trained MLP layers’ mappings. One compelling alternative to MxDs for efficient conditional computation is the $\mu$MoE layer [40], which imposes low-rankness on expert weights to achieve parameter-efficiency. Whilst $\mu$MoEs are found to perform competitively in the pre-training setting, the impact of low-rankness on approximations of existing layers will determine their suitability in the sparse layer approximation setting studied in this work.
@@ B.4.1 Comparisons to low-rank MoEs
We match the MxD experimental configuration as closely as possible for a fair comparison. For the encoders, we mirror MxDs and use the GELU activation function, which we find through ablations in Section B.6 to perform the best. We initialize the parameters the same as MxDs and Skip Transcoders: we use the standard PyTorch linear layer initialization for $\mathbf{D},\,\mathbf{C}$ (and the encoder layers), and initialize $\mathbf{W}$ as the zero matrix.
Figure 12: Comparisons to $\mu$MoEs for various choices of (normalized) rank: high rank weights best-preserve the models’ downstream cross-entropy loss.
We vary the $\mu$MoE layer rank $R$, training fully 3 sparse approximation layers for $K=32$ active experts, varying the total number of experts $N$ to keep the parameter count the same–isolating the impact of the choice of rank. As with the main experiments, we record the downstream model loss when we splice in the trained layer to replace the MLP layers, shown in Figure 12.
As can be seen, the $\mu$MoE layers perform well when they are close to full-rank (i.e. when the normalized rank $\frac{R}{O}\rightarrow 1$). Crucially, however, performance drops off notably when the rank is reduced. Whilst $\mu$MoEs still perform far better than neuron-level sparsity methods (i.e. the corresponding CE loss results in Figure 3), we observe that full-rankness is necessary for the most faithful layer approximations–which the proposed MxDs provide by design.
As a motivating example, for why SparseMoEs and SoftMoEs are not practical: SparseMoEs [36] and SoftMoEs [70] require $2.16$ trillion parameters for a single layer, for the same $86$k experts we use for Llama-3.2-3B. This is orders of magnitude more parameters than the entire base network itself, making it prohibitively costly for SparseMoEs to scale to sufficiently high expert counts.
B.4.2 MxD empirical expert rank
Next, we show experimentally that the learned experts’ matrices $\mathbf{W}_{n}=\mathbf{D}\,\text{diag}({\mathbf{c}}_{n})\in\mathbb{R}^{H\times O}$ are very nearly full-rank in practice, corroborating the properties of expert matrices shown theoretically in Lemma 1. We compute the mean ‘normalized rank’, which we take for MxDs to be the empirical matrix rank of the learned expert’s weights, over the maximum possible rank given the dimensions:
@@ B.4.2 MxD empirical expert rank
We show in Table 3 the normalized rank across all 4 base models: MxD’s learned experts exhibit no rank deficiencies, providing further evidence of the large potential capacity of MxD layers despite their sparsity constraints on the expert-level.
Table 3: Mean normalized expert matrix rank of Equation 23 across models for the first 2k experts in $K=32$ trained MxDs – the learned expert matrices are very close to full column rank.
@@
$0.99\pm 0.005$ $0.99\pm 0.007$ $0.99\pm 0.005$ $0.99\pm 0.002$
B.5 Sparse probing

### v2

@@ A.1 Proof of rank equality
Equation˜6 shows that for a fixed expert $n$, the $n^{\text{th}}$ row $\mathbf{c}_{n}\in\mathbb{R}^{O}$ essentially scales the columns of matrix $\mathbf{D}\in\mathbb{R}^{H\times O}$. This is equivalent to right-multiplying matrix $\mathbf{D}$ by a diagonal matrix formed from $\mathbf{c}_{n}\in\mathbb{R}^{O}$. Indeed, the $(h,o)$ entry of such matrix product is
@@
since all off-diagonal terms (i.e., $i\neq o$) in Equation˜7 vanish and $\text{diag}(\mathbf{c}_{n})_{oo}=c_{no}$ by construction. Comparing Equation˜6 and Equation˜9 shows that, for every $h\in\{1,2,\ldots,H\}$ and $o\in\{1,2,\ldots,O\}$ we have
@@
Finally, a standard result in linear algebra [97] has that $\text{rank}(\mathbf{AB})=\text{rank}(\mathbf{A})$ for any $\mathbf{A}\in\mathbb{R}^{H\times O}$ and invertible matrix $\mathbf{B}\in\mathbb{R}^{O\times O}$. Since matrix $\mathrm{diag}\left({\mathbf{c}}_{n}\right)\in\mathbb{R}^{O\times O}$ is invertible by assumption in Lemma˜1, setting $\mathbf{A}=\mathbf{D}$ and $\mathbf{B}=\text{diag}({\mathbf{c}}_{n})$ yields the rank equality. ∎
@@ A.2 Proof of MxD forward pass equivalence
Recall we have input vector $\mathbf{z}\in\mathbb{R}^{H}$, expert coefficients $\mathbf{a}\in\mathbb{R}^{N}$, and layer weights $\boldsymbol{\mathcal{W}}\in\mathbb{R}^{N\times H\times O}$. The weights are defined in Equation˜4 element-wise through the Hadamard product * as
@@
for learnable parameters $\mathbf{C}\in\mathbb{R}^{N\times O},\,\mathbf{D}\in\mathbb{R}^{H\times O}$. Lemma˜2 states that MxD’s forward pass can be equivalently expressed as
@@
Plugging in the definition of $\mathbf{w}_{nh:}\in\mathbb{R}^{O}$ from Equation˜4 then yields
@@
which is exactly the RHS of Equation˜5, showing the MxD forward pass is equivalent to the Hadamard product of $\mathbf{C}^{\top}\mathbf{a}$ and $\mathbf{D}^{\top}\mathbf{z}$. ∎
@@ A.3 Intuition for weight parameterization through the lens of tensor methods
A second complementary way of viewing the MxD layer’s parameterization (and its full-rank properties) is through the lens of tensor methods [34]. A tensor-based motivation for MxD’s weight tensor parameterization and forward pass is presented in Section˜A.3.1 and Section˜A.3.2, respectively.
@@ A.3.1 MxD weight tensors through the Khatri-Rao product
MxDs construct the collective weight tensor through the Khatri-Rao product $\odot$ [34] of the two factor matrices $\mathbf{C}\in\mathbb{R}^{N\times O},\,\mathbf{D}\in\mathbb{R}^{H\times O}$. Concretely, the mode-$3$ unfolding222which is simply a reshaping of a higher-order tensor into a matrix, arranging all $N$ expert matrices’ column vectors along the columns of a new matrix. of the third-order weight tensor $\boldsymbol{\mathcal{W}}\in\mathbb{R}^{N\times H\times O}$ in MxDs from Equation˜4 is alternatively given by:
@@ A.3.2 Tensorized MxD forward pass
Furthermore, the layer’s forward pass can then be viewed as performing two tensor contractions between the third-order weight tensor $\boldsymbol{\mathcal{W}}\in\mathbb{R}^{N\times H\times O}$ (collecting all $N$ experts’ $H\times O$-dimensional matrices) and expert coefficients $\mathbf{a}\in\mathbb{R}^{N}$ and hidden activations $\mathbf{z}\in\mathbb{R}^{H}$. This can be expressed in terms of the so-called mode-$n$ product (denoted by $\times_{n}$) [34] as follows:
@@ A.4 GLU encoders are a mixture of rank-1 linear experts
for input vector $\mathbf{x}\in\mathbb{R}^{I}$, learnable weights $\mathbf{E}_{\text{GLU}},\,\mathbf{E}\in\mathbb{R}^{I\times H}$, and activation function $\psi(.)$. To transform Equation˜17 into the same model form as MxDs, we first pre-multiply the LHS by the identity matrix to match the MxD model form of Equation˜5, yielding:
@@
where $(\mathbb{I})_{n}\in\mathbb{R}^{H}$ is the $n^{\text{th}}$ row of the $H$-dimensional identity matrix (i.e. a one-hot vector with its only non-zero element at index $n$). We draw particular attention to how the $n^{\text{th}}$ expert’s matrix $\mathbf{W}_{n}=\mathbf{E}\,\mathrm{diag}\left((\mathbb{I})_{n}\right)\in\mathbb{R}^{I\times H}$ essentially picks out the $n^{\text{th}}$ column of $\mathbf{E}$, leaving all remaining H1H-1 columns as zero vectors. Therefore, GLU encoders compute a MoE with linear expert weights of (at most) rank 1. This relationship between GLUs and conditional computation is consistent with prior work interpreting individual GLU column vectors as experts [99]. Whilst GLUs’ encoders’ model form does not put any inherent restrictions on the total number of rank-1 terms that can contribute to the output, the sparsity necessary for specialization does.
@@
    Replace $\mathbb{I}$ in Equation˜18 with learnable, non-diagonal weight matrices for full-rankness.
@@ A.5 Hadamard-factorized tensors generalize MoVs
Prior work [37] proposes to linearly combine $N$ many (IA)3 adapters [38] for parameter-efficient MoEs for instruction fine-tuning. The implementation results in a very similar functional form to the factorized forward-pass in MxDs. Interestingly, the Hadamard product parameterization of the third-order weight tensor in Equation˜4 provides a more general framework through which one can also derive MoVs’ model form, shedding light on the relationship to the proposed MxDs and their benefits. Concretely, factorizing the weight tensor instead along the second mode as $\boldsymbol{\mathcal{W}}(n,:,o)=\mathbf{c}_{n}*\mathbf{d}_{o}\in\mathbb{R}^{H}$ in our framework immediately recovers MoV [37] as a special case. In particular, in contrast to the MxD in Section˜A.3 whose weight tensor can be parametrized equivalently through its mode-$3$ unfolding [34], MoV’s implicit weight tensor can be given in terms of its mode-$2$ unfolding in terms of a similar Khatri-Rao product of two factor matrices.
Instead, MoVs in analogy would yield expert weights by pre-multiplying $\mathbf{D}$ as: $\mathbf{W}_{n}=\mathrm{diag}(\mathbf{c}_{n})\,\mathbf{D}\in\mathbb{R}^{H\times O}$ for much larger $\mathbf{C}\in\mathbb{R}^{N\times H}$). Due to $H\gg O$, our proposed MxD formulation yields around $4\times$ the number of specialized units as MoVs with the same parameter budget (yet MoVs’ experts are of no higher rank than MxDs’), making MxDs a much more suitable and efficient class of layer for our goal of scalable specialization. We therefore see that the proposed lens of tensor methods for unification provides valuable insights about how to design more interpretable layers with the minimum trade-off to capabilities.
A.6 TopK activation function
Formally, for activations $\mathbf{z}\in\mathbb{R}^{H}$ (possibly with the ReLU activation applied), the TopK activation function can be formulated elementwise as the following:
$\text{TopK}(\mathbf{z})_{h}=\begin{cases}z_{h},&\text{if }z_{h}\geq\tau_{K}(\mathbf{z})\\ 0,&\text{otherwise}\end{cases},$
where $\tau_{K}(\mathbf{z})\in\mathbb{R}$ returns the value of the $K^{\text{th}}$ largest element of vector $\mathbf{z}\in\mathbb{R}^{H}$. Intuitively, the TopK activation retains only the largest $K\leq H$ elements of the $H$ neurons, setting the rest to zero. We note here that the definition above does not handle ties (although this is very unlikely in practice, given the activations are continuous).
@@ B.1 Faithfulness in output space
Experiments in Section˜3.1.2 of the main paper report the number of future predicted words that are identical from the original model and from the LLM with the sparse layer replacements. We also show qualitative examples of the first $8$ prompts and the subsequent ‘diffs’ (using Python 3’s difflib) of the generated tokens in Figures˜7 and 8–we see MxDs’ superior ability to preserve model functionality as it propagates through to the output space of future tokens.
@@ B.2 Additional reconstruction metrics
To highlight the scale of difference in the reconstructions between MxDs and the baselines, we also plot in Figure˜9 the normalized MSE at the end of training for all models and LLMs. At the smallest values of $K$ (which we care about most for interpretability), MxDs’ normalized MSE is up to an order of magnitude smaller than Transcoders’.
@@ B.3 Results on additional layers
We also fully train all models and baselines (with 4 different values of $K$) on different target layers for each model. The results are shown in Figure˜10 for $48$ additional trained layers for the same setup in the original paper, using different colours to highlight that these are new results. As can be seen, the same trend holds: MxDs significantly outperform the baselines at small $K$ in all LLMs.
@@
B.4 Initial results on Gemma2-27B
In the main paper, we perform experiments on base models with up to 3B parameters. For initial evidence of scalability to even larger models, we perform here additional large-scale experiments on the 27B Gemma2 model (in half-precision, and with a smaller batch size to fit into memory). We see MxDs continue to outperform on the sparsity-accuracy frontier, as shown by the (normalized) reconstruction loss in Table˜3.
Table 3: Normalized MSE ($\downarrow$) on Gemma2-27B, Layer 20, $K=32$, partially trained for 50k iterations. To fit in memory, the model is loaded in half precision. We use $1/8$ the batch size and # buffers stored, and $1/2$ the multiplier on the number of latent features (with values of 4, 16, and 16, respectively).
Iterations 10k 20k 30k 40k 50k
Transcoder
0.147 0.132 0.128 0.123 0.119
Skip Transcoder
0.128 0.107 0.102 0.100 0.093
MxD
0.098 0.096 0.093 0.086 0.069
B.5 Expert rank
This section concerns the matrix rank of the linear experts in parameter-efficient MoEs. We first compare to low-rank MoEs in Section˜B.5.1 to demonstrate the benefits of full-rankness, and then follow up in Section˜B.5.2 by confirming that the learned MxD expert ranks are close to maximum in the trained models.
B.5.1 Comparisons to low-rank MoEs
[…]
@@ B.5.1 Comparisons to low-rank MoEs
We match the MxD experimental configuration as closely as possible for a fair comparison. For the encoders, we mirror MxDs and use the GELU activation function, which we find through ablations in Section˜B.8 to perform the best. We initialize the parameters the same as MxDs and Skip Transcoders: we use the standard PyTorch linear layer initialization for $\mathbf{D},\,\mathbf{C}$ (and the encoder layers), and initialize $\mathbf{W}$ as the zero matrix.
[…]
As a motivating example, for why SparseMoEs and SoftMoEs are not practical: SparseMoEs [36] and SoftMoEs [70] require $2.16$ trillion parameters for a single layer, for the same 8686k experts we use for Llama-3.2-3B. This is orders of magnitude more parameters than the entire base network itself, making it prohibitively costly for SparseMoEs to scale to sufficiently high expert counts.
B.5.2 MxD empirical expert rank
Next, we show experimentally that the learned experts’ matrices $\mathbf{W}_{n}=\mathbf{D}\,\text{diag}({\mathbf{c}}_{n})\in\mathbb{R}^{H\times O}$ are very nearly full-rank in practice, corroborating the properties of expert matrices shown theoretically in Lemma˜1. We compute the mean ‘normalized rank’, which we take for MxDs to be the empirical matrix rank of the learned expert’s weights, over the maximum possible rank given the dimensions:
@@ B.5.2 MxD empirical expert rank
We show in Table˜4 the normalized rank across all 4 base models: MxD’s learned experts exhibit no rank deficiencies, providing further evidence of the large potential capacity of MxD layers despite their sparsity constraints on the expert-level.
Table 4: Mean normalized expert matrix rank of Equation˜23 across models for the first 2k experts in $K=32$ trained MxDs – the learned expert matrices are very close to full column rank.
@@
$9\pm 0.005 0.99±0$$9\pm 0.007$ $9\pm 0.005 0.99±0$$9\pm 0.002$
B.6 Computational cost of sparse layers
We show here that there is minimal difference between the computational cost of sparse layer variants. We first report theoretical layer FLOPs, and then report empirical benchmarks.
Parameter count & FLOPs
We first tabulate in Table˜5 the theoretical parameter counts and inference-time FLOPs for MxDs vs Sparse MLPs. To be consistent with the popular PyTorch library [102] we count one fused multiply-add as one FLOP, and count the Hadamard product between two $d$-dimensional vectors as requiring $d/2$ FLOPs. Note that, For a chosen expert count $N$, we set the width of Sparse MLPs (e.g. Transcoders) to $H:=N+H^{*}$ to parameter-match the models.
Table 5: Here $I,O$ denotes the input and output dimensions, $H^{*}$ the original model’s hidden width, NN the MxD expert count, and $H$ the width of the sparse MLPs.
Parameter count FLOPs
Sparse MLP H(I+O)H\,(I+O) H(I+O)H\,(I+O)
MxD $N\,(I+O)+H^{*}\,(I+O)$ $N\,(I+O)+H^{*}\,(I+O)+O/2$
Empirical benchmarks
We next run benchmarks for a Sparse MLP layer vs MxD with a batch size of 512, and dimensions: $I=H^{*}=O=1024$, and number of experts/features as N=8192,H=9216N=8192,\,H=9216. The results are tabulated in Table˜6, where we see performance (and cost) is similar for the two layers.
Table 6: Here $I,O$ denotes the input and output dimensions, $H^{*}$ the original model’s hidden width, NN the MxD expert count, and $H$ the width of the sparse MLPs.
Peak memory usage (MiB) Latency (ms) Parameter count Reported FLOPs ([102])
Sparse MLP 386.50 1.394 18,874,368 18,874,368
MxD 389.50 1.457 18,874,368 18,874,880
B.7 Sparse probing